identifying the unknown using IR, HNMR CNMR and Mass spectrums

Unknown H407 100- 18-N-9500 Relative Intensity 160 180 200 220 Order 20 40 60 80 100 120 140 m/z

Unknown H407 mm NREUTHER

Unknown H407 d. 3H 1 Зн m, 3H d. 1H m, 2H m, 2H m, 1Η br. s, 1H 10 8 PPM

Unknown H407 20 0 80 60 4 0 100 120 140 180 160
Answer #1

|Answer A] Mass Speckum Data m12= 192(M+.), 147(100), 118,92 aj The mass of molecular i on peak at m2 = 192, indicate that moD] +H-NMR Data Sppm 0.90(3H, t] =) Cez che 1.0 (3H, d) = CH-CH- 1.7(2H, M) = -cthy 2.3(1H, m] = CH- 3.8(1H1, d) = -9 CH CH 7.on the basis of IR, NMR, emr and mass, the Possible structure of Compound is Propel.corbes 2-3cm) o choch CH CH 0.90lt) W e a